OpenStudy (anonymous):

Find the exact value of sin((cos^-1)(5/13))

OpenStudy (anonymous):

draw a right triangle and label the sides according to what's given... let \(\large \theta=cos^{-1}(\frac{5}{13}) \) now, \(\large \theta=cos^{-1}(\frac{5}{13}) \) means \(\large cos\theta=\frac{5}{13} \) so you can draw the right triangle along with the angle theta...